# Cost-Effective Online Multi-LLM Selection with Versatile Reward Models

### v1

@@ 1 Introduction
In today’s digital era, Large Language Models (LLMs) like ChatGPT and Claude lead innovations in computational linguistics and cognitive processing, revolutionizing digital interaction through advancements in text generation, sentiment analysis, translation, and comprehension [61, 60]. The emergence of numerous high-performance LLMs has sparked significant interest in the challenge of model selection [19, 20, 64]. Typically, current schemes for selecting LLMs often rely on identifying the best-performing model under a static setting, such as selecting the LLM that achieves the lowest perplexity score [53, 49]. Yet, the diverse capabilities of various LLMs present an opportunity to adopt a task-specific selection approach, where different LLMs have their own strengths and weaknesses, e.g., Investlm [67] is specifically designed for the financial sector and may better suit queries related to investments. Moreover, the limitations of static selection methods become more pronounced due to factors like “generation diversity”, where a less expensive LLM may perform better in certain scenarios [12], and “data drift”, which refers to changes in the characteristics of answers generated in real-time compared to those in the training data [5]. These issues underscore the importance of developing a dynamic, online approach for selecting the “appropriate” LLMs for the varying performance levels of different models and various application needs.
Furthermore, scenarios combining multiple LLMs (or agents) to complete tasks have commonly emerged, moving beyond using a single LLM. Platforms like Poe [50] have spearheaded functionalities that integrate several bots within a single chat session. [42] introduces a dynamic LLM-agent network through dynamic interaction architecture and intelligent agent team optimization. [28] proposes a meta-programming framework that enhances the collaboration of multiple LLMs. [24] explores the implementation of LLM cascades for generative tasks. However, previous works have not considered optimizations tailored to the characteristics of different tasks, which may feature different forms of rewards. Below, we present three streamlined collaborative combinations of LLMs for different tasks: 1. For user experience enhancement, multiple LLMs may be deployed to ensure satisfactory outcomes. 2. In educational tutoring, subject-specific LLMs operate in parallel, with failures in one not severely impacting others. 3. In project development, LLMs manage different sub-modules, where any module’s failure could jeopardize the entire project. These three examples underscore the importance of combining multiple suitable LLMs based on task structure.
@@
Based on the discussions above, we introduce the Cost-effective Combinatorial Multi-armed Bandit with Versatile reward models (C2MAB-V), designed to synergize the integration of diverse LLMs across different task types. C2MAB-V manages the dual challenges of selecting LLMs that both achieve high performance and meet cost constraints. Additionally, C2MAB-V adapts to various multi-LLM collaborative tasks by utilizing a “combinatorial model selection strategy” which extends beyond traditional single-model limitations by encompassing a broad spectrum of LLM candidates. Finally, to address the NP-hard complexities of combinatorial LLM selection under cost considerations, we transform the initial problem into a continuous form. This process is executed on the local server, which is required to handle user queries. Meanwhile, our discretization method, which guarantees precise evaluation of LLM performance post-relaxation, is implemented on the scheduling cloud where multiple LLMs are deployed. This two-tiered approach can accommodate both the limited resources of the local server and strategically mitigate the scheduling cloud’s workload.
@@
Novel Multi-LLM Selection Formulation. We introduce a novel formulation of the cost-effective multi-LLM selection, designed for tasks requiring collaboration among multiple LLMs with varying reward structures. This formulation emphasizes the essential balance between exploring new models and exploiting proven effective models, while adhering to long-term cost budget considerations and securing high rewards across diverse multi-LLM tasks.
@@ 2.1 Related Work
Multi-LLM Combination. The combination of multiple LLM models for enhanced performance has received considerable attention, aimed at bolstering output quality [31]. Techniques such as “knowledge distillation[25, 54] facilitate the training of compact models to mimic the behavior of larger, more complex models, thereby optimizing resource utilization. Meanwhile, “ensemble learning[27, 29] leverages the collective predictions of independently trained models for superior results. Nonetheless, the prevalent practice of withholding model internals from commercial LLM, such as OpenAI and Claude [47, 14], constrains knowledge distillation by obscuring the “teacher” model’s intricacies, which may diminish the “student” model’s replication efficacy. Concurrently, ensemble learning faces increased financial and procedural burdens due to the necessity of amalgamating diverse model outputs, a task complicated by the absence of open-source models. Our approach distinctively accounts for the online aspects of model selection, the diversity of multi-LLM task types, and the associated costs, setting it apart from previous methods.
@@ 3 Problem Formulation
In this section, we introduce our online framework of a cost-effective combinatorial bandit for multi-LLM selection with versatile reward models, namely, C2MAB-V. (Refer to Fig. 3 for details.)
Local-Cloud Architecture. Given the large number of parameters and significant storage overhead of LLMs, a typical approach, if not opting for a streamlined but less capable version of LLMs, involves deploying multiple LLMs on a resource-abundant scheduling cloud. When responding to user queries, the local server first handles the requests and synchronizes communication with the cloud to initiate tasks (for discussions on asynchronous handling, see Appendix C.2). Subsequently, user feedback is stored on the local server. [^1] Feedback includes both direct user input and data from techniques that quantify user behavior or responses, as well as preference simulators [17]. Regardless of the methods, we refer to all these collectively as “feedback.” While the scheduling cloud serves multiple local servers, for ease of presentation, we focus on describing the relationship between a single local server and the cloud.
@@
Online Learning Protocol. A combinatorial LLM instance involves the sequential interaction between the local server, the scheduling cloud, and user queries within our local-cloud architecture. The local server processes user activity and feedback, particularly focusing on the users’ performance feedback regarding the utilized LLMs. This process involves locally recording and updating the performance evaluations of the LLM. Then, the local server will transmit these information to the scheduling cloud. The scheduling cloud undertakes the new coordination and selection of an action $S_{t}$ (i.e., selecting a subset of LLMs for service) after receiving the newly updated information from the local server. Such performance evaluations, termed as rewards, could, for instance, be based on the ROUGE-2 score for automatic summarization tasks [37], and are represented by a random variable vector $\boldsymbol{X}_{t}=(X_{t,1},\ldots,X_{t,|S_{t}|})$. For simplification, we posit that $\boldsymbol{X}_{t}\in[0,1]^{|S_{t}|}$. Furthermore, for each LLM $k\in\mathcal{K}$ included in the action $S_{t}$, its associated cost $y_{t,k}$ is observable at round $t$, with its time-invariant expected value denoted by $c_{k}\coloneqq\mathbb{E}[y_{t,k}]$.
Versatile Reward Models. Let $\boldsymbol{\mu}=(\mu_{1},...,\mu_{K})$ represent the “initially unknown” mean vector of outcomes for each LLM with $\mu_{k}=\mathbb{E}[X_{t,k}]$. We consider versatile reward models for different combinations of LLMs in multi-LLM tasks below. For more discussions, please refer to Section A.1.
@@
  • All In Combination (AIC): $r(S;\boldsymbol{\mu})=\prod_{k\in S}\mu_{k}.$ This reward model is exemplified in development tasks, where each LLM is responsible for developing sub-modules. The key aspect of this reward model is that the failure of any LLM leads to the failure of the entire development task, thus ensuring the success of the whole collaborative work.
Partial LLM Feedback. In the process of selecting and querying LLM in action $S_{t}$, not all LLMs may actually be queried in the muti-LLM tasks. Consequently, the local server often only observes outcomes from a subset of the selected LLMs. Formally, this means the local server observes feedback from the first $\mathcal{F}_{t}$ LLM in action $S_{t}$, where its cardinality $F_{t}=\min\{k\in S_{t}:X_{t,k}=1\}$ and $F_{t}=|S_{t}|$ if all LLMs are queried in the selected action $S_{t}$.
Budget Cost Consideration. The total cost associated with the action $S_{t}$, chosen by the scheduling cloud at round $t$, is quantified as $\sum_{k\in S_{t}}y_{t,k}$, where this cost is generally proportional to the query length, determined by the token count in queries [12]. Additionally, a predetermined budget guarantee threshold $\rho>0$ exists for the combinatorial set of LLM, requiring the cumulative cost of the selected action to remain beneath this threshold $\rho$ over the long term. This mechanism allows organizations, including enterprises and governmental bodies, to manage LLM usage efficiently and align with budgetary constraints. To assess compliance with this budgetary constraint, the concept of constraint violation, following [7], is introduced. For $T>0$, the violation is defined by:
$V(T)=\left[\frac{1}{T}\sum_{t=1}^{T}\sum_{k\in S_{t}}y_{t,k}-\rho\right]^{+},$ (1)
where $[x]^{+}=\max(x,0)$ delineates the non-negative portion of $x$. Although it’s conceivable to occasionally exceed this cost constraint during the online learning of LLM performance phase, significant overruns are not permissible, thereby necessitating the violation metric. A violation rate decreasing as $\tilde{O}\left(T^{-\gamma}\right)$ signifies effective constraint adherence, indicating that $V(T)$ diminishes at this rate, where $\gamma>0$ illustrates the rate of reduction in violations with increasing $T$ [43, 11].
@@
Goal of Multi-LLM Selection. The performance of an online algorithm $A$ is critically evaluated through both violation and regret metrics, as defined in Eq. (1) and Eq. (2), respectively. A diminishing regret implies an increasingly oracle-like performance by algorithm $A$, indicative of its efficiency in selecting optimal LLM combinations. Conversely, a reduction in violation highlights improved compliance with predefined cost constraints over time. The dual objectives for algorithm $A$ involve simultaneously minimizing regret and violation under different reward models. [^3] Without loss of generality, we assume $\mu_{k},c_{k}\in[0,1]$ for all $k\in\mathcal{K}$.
@@ 4 Algorithm Design
Figure 3: Design of online C2MAB-V flow.
Next, we present the design of the online C2MAB-V framework, as depicted in Algorithm 1. Our method addresses the challenge of estimating unknown rewards and costs associated with using corresponding LLMs. As shown in Fig. 3, given the NP-hardness complexity of selecting a combination of LLMs under cost constraints, we transform the original integer problem into a continuous space. Utilizing the limited resources of a local server, we solve this relaxed optimization problem, which can also ease the computational load on the scheduling cloud when supporting multiple local-server query requests. A scheduling cloud with multiple LLMs coordinates and selects based on continuous data transmitted from the local server. Meanwhile, the local server collects user feedback to enhance the online evaluations of LLMs. Additionally, the cloud does not have access to the original sensitive user data, enhancing future privacy protection prospects. Next, we describe the processes undertaken by the local server and the scheduling cloud, respectively.
@@
1:Set of all LLMs $\mathcal{K}$, cost constraint $\rho$, $\delta\in(0,1]$.
@@
4:     Predict reward of $k$-th LLM $\bar{\mu}_{t,k}=\min\{\hat{\mu}_{t,k}+\alpha_{\mu}\rho_{t,\mu_{k}},1\}$;
5:     Estimate cost of $k$-th LLM $\underaccent{\bar}{c}_{t,k}=\max\{\hat{c}_{t,k}-\alpha_{c}\rho_{t,c_{k}},0\}$;
@@
8:     Observe the corresponding reward of LLMs for $k\in\mathcal{F}_{t}$ and the cost of LLM in $S_{t}$;
@@ 4.1 Procedures by Local Server
Confidence Bound for Reward and Cost. To avoid the limitations of a greedy LLM selection strategy, i.e., $\operatorname*{arg\,max}_{k\in\mathcal{K}}$ $\hat{\mu}_{t,k}$ with $\hat{\mu}_{t,k}$ denoting the empirical estimate of the true mean of LLM $k$, which might result in the risk of overlooking superior LLM options, we implement the confidence bound (CB) method [35, 41, 65]. This approach promotes exploration among various LLM alternatives, thereby reducing the likelihood of consistently choosing sub-optimal options without considering potentially better LLMs. We define the confidence radius of CB as $\rho_{t,\mu_{k}}=\sqrt{\ln\left(\frac{2\pi^{2}Kt^{3}}{3\delta}\right)/2T_{t,\mu_{k}}}$, which quantifies the exploration potential for the reward of LLM $k$ at round $t$, with $\delta$ in the range $(0,1]$. The term $T_{t,\mu_{k}}$ represents the number of times LLM $k$ has been selected in action $S_{t}$. Accordingly, the adjusted reward prediction for LLM $k$ is defined as $\bar{\mu}_{t,k}=\min\{\hat{\mu}_{t,k}+\alpha_{\mu}\rho_{t,\mu_{k}},1\}$, with $\alpha_{\mu}$ denoted as a positive control parameter. Furthermore, we take into account the cost associated with each LLM. In light of the uncertainty introduced by complex queries, we adopt a cautious approach to ensure adherence to the cost consideration. Specifically, the adjusted cost estimate for LLM $k$ at round $t$, $\underaccent{\bar}{c}_{t,k}$, is determined by $\underaccent{\bar}{c}_{t,k}=\max\{\hat{c}_{t,k}-\alpha_{c}\rho_{t,c_{k}},0\}$. Here, $\hat{c}_{t,k}$ is the empirical cost, $\alpha_{c}$ is a positive adjustment parameter, and $\rho_{t,c_{k}}=\sqrt{\ln\left(\frac{2\pi^{2}Kt^{3}}{3\delta}\right)/2T_{t,c_{k}}}$, with $T_{t,c_{k}}$ denoting the count of LLM $k$ selection up to round $t$. Using this CB approach, we have the following lemma.
@@
Please refer to Appendix B.1 for the proof. Lemma 1 underscores the high-probability events wherein the empirical estimates for both reward and cost closely align with their true means of LLMs.
@@
    Greedy algorithms, apt for such constrained continuous problems, can efficiently select $\tilde{z}_{k}$ values that optimize $1-\bar{\mu}_{k}\tilde{z}_{k}$ within constraints [6]. Subsequently, this optimization problem can be easily and effectively solved using a common greedy algorithm.
@@
The commonalities in the above three types of reward forms are in Appendix A.2. We will further describe the selection of multiple LLMs based on the continuous variable $\tilde{\boldsymbol{Z}}_{t}=\{\tilde{z}_{t,1},\tilde{z}_{t,2},\cdots,\tilde{z}_{t,K}\}\in[0,1]^{K}$ at round $t$ in Section 4.2 on Procedures by Scheduling Cloud.
@@
$\hat{\mu}_{t+1,k}=\cfrac{T_{t,\mu_{k}}\hat{\mu}_{t,k}+X_{t,k}}{T_{t+1,\mu_{k}}},\quad k\in\mathcal{F}_{t},\quad\hat{c}_{t+1,k}=\cfrac{T_{t,c_{k}}\hat{c}_{t,k}+y_{t,k}}{T_{t+1,c_{k}}},\quad k\in S_{t}.$ (6)
Note that for an action $S_{t}$, the local server only monitors the performance of the LLMs that are actually used in $S_{t}$, as not all LLMs are utilized for every task type. For example, in the case of an AWC task type, if one LLM provides a satisfactory answer, the remaining LLMs are not utilized. On the other hand, the cost for all LLMs involved in an action $S_{t}$ can be calculated based on the tokens and price criteria of each LLM. This accounts for the difference in the range setting of $k$ in Eq. (6).

### v2

@@ 1 Introduction
In the digital era of today, large language models (LLMs) such as ChatGPT lead innovations in computational linguistics and cognitive processing [65]. The emergence of numerous high-performance LLMs has sparked significant interest in the challenge of model selection [22, 68]. Typically, current schemes for selecting LLMs often rely on identifying the best-performing model in a static setting, such as selecting the LLM that achieves the lowest perplexity score [57, 53]. However, the diverse capabilities of various LLMs present an opportunity to adopt a task-specific selection approach, where different LLMs have their own strengths and weaknesses, e.g., Investlm [71] is specifically designed for the financial sector and may better suit queries related to investments. Moreover, the limitations of static selection methods become more pronounced due to factors like “generation diversity”, where a less expensive LLM may perform better in certain scenarios [13], and “data drift”, which refers to changes in the characteristics of answers generated in real-time compared to those of training data [6]. Consequently, beyond large-scale pre-training methods, these issues highlight the need for an “online” approach to optimize decision-making in selecting the “appropriate” LLMs, which leverage continuous feedback to adapt to the varying performance levels of models and diverse application needs through ongoing user interaction.
Furthermore, scenarios combining multiple LLMs (or agents) to complete tasks have commonly emerged, moving beyond using a single LLM. Platforms like [54] have spearheaded functionalities that integrate several bots within a single chat session. [46] introduces a dynamic LLM-agent network through dynamic interaction architecture and intelligent agent team optimization. [31] proposes a meta-programming framework that enhances the collaboration of multiple LLMs. [26] explores the implementation of LLM cascades for generative tasks. However, previous works have not considered optimizations tailored to the characteristics of different tasks, which may feature different forms of rewards. Below, we present three streamlined collaborative combinations of LLMs for different tasks: 1. For user experience enhancement, multiple LLMs may be deployed to ensure satisfactory outcomes. 2. In educational tutoring, subject-specific LLMs operate in parallel, with failures in one not severely impacting others. 3. In project development, LLMs manage different sub-modules, where any module’s failure could jeopardize the entire project. These three examples underscore the importance of combining multiple suitable LLMs based on the specific “task structure”.
@@
Based on the discussions above, we introduce the Cost-effective Combinatorial Multi-armed Bandit with Versatile reward models (C2MAB-V), designed to synergize the integration of diverse LLMs across different task types. C2MAB-V manages the dual challenges of selecting LLMs that both achieve high performance and meet cost constraints. Additionally, C2MAB-V adapts to various multi-LLM collaborative tasks by utilizing a “combinatorial model selection strategy” which extends beyond traditional single-model limitations by encompassing a broad spectrum of LLM candidates. Finally, to address the NP-hard complexities of combinatorial LLM selection under cost considerations, we transform the initial problem into a continuous form. This process is executed on the local server, which is required to handle user queries. Meanwhile, our discretization method, which guarantees precise evaluation of LLM performance post-relaxation, is implemented on the scheduling cloud where multiple LLMs are deployed. This two-tiered approach can accommodate both the limited resources of the local server and strategically mitigate the scheduling cloud’s workload. We emphasize that this paper focuses on a fully online learning approach that operates without initial information and relies solely on user feedback with robust interpretation. Our method is designed to complement, not compete with, current approaches that utilize GPU resources for offline training, where offline training can streamline the search space and enhance the convergence speed of online learning.
@@
Novel Multi-LLM Selection Formulation. We introduce a novel formulation of the cost-effective multi-LLM selection, designed for tasks requiring collaboration among multiple LLMs with versatile reward structures. This formulation emphasizes the essential balance between exploring new models and exploiting proven effective models, while adhering to long-term cost budget considerations and securing high rewards across diverse multi-LLM tasks.
@@ 2.1 Related Work
Multi-LLM Combination. The combination of multiple LLM models for enhanced performance has received considerable attention, aimed at bolstering output quality [35]. Techniques such as “knowledge distillation[28, 58] facilitate the training of compact models to mimic the behavior of larger and more complex models, thus optimizing resource utilization. Meanwhile, “ensemble learning[30, 32] leverages the collective predictions of independently trained models for superior results. Nonetheless, the prevalent practice of withholding model internals from commercial LLM, such as [52], restricts knowledge distillation by obscuring the “teacher” model, which can decrease the replication efficacy of the “student” model. Concurrently, ensemble learning faces increased financial and procedural burdens due to the necessity of amalgamating various model outputs, a task complicated by the absence of open-source models. [47, 18, 13] primarily address one of the scenarios we have outlined, i.e., selecting a single LLM that best satisfies the user from multiple options. Moreover, these approaches do not incorporate user feedback. Our approach distinctively accounts for the online aspects of model selection, the diversity of multi-LLM task types, and the associated costs, setting it apart from previous methods.
@@ 3 Problem Formulation
In this section, we introduce our online framework of a cost-effective combinatorial bandit for multi-LLM selection with versatile reward models, namely, C2MAB-V (Please refer to Fig. 3 for details). A summary of the main symbols is provided in Appendix A.
Local-Cloud Architecture. Given the large number of parameters and significant storage overhead of LLMs, a typical approach, if not opting for a streamlined but less capable version of LLMs, involves deploying multiple LLMs on a resource-abundant scheduling cloud. When responding to user queries, the local server first handles the requests and synchronizes communication with the cloud to initiate tasks (for discussions on asynchronous handling, see Appendix E.3). Subsequently, user feedback is stored on the local server. [^1] Feedback can include both direct user input and data from techniques that quantify user behavior or responses, as well as preference simulators [20]. Regardless of the methods, we refer to all these collectively as “feedback” for simplicity. While the scheduling cloud serves multiple local servers, for ease of presentation, we focus on describing the relationship between a single local server and the cloud.
@@
Online Learning Protocol. A combinatorial LLM instance involves the sequential interaction between the local server, the scheduling cloud, and user queries within our local-cloud architecture. The local server processes user activity and feedback, particularly focusing on the users’ performance feedback regarding the utilized LLMs. This process involves locally recording and updating the performance evaluations of the LLM. Then, the local server will transmit these information to the scheduling cloud. The scheduling cloud undertakes the new coordination and selection of an action $S_{t}$ (i.e., selecting a subset of LLMs for service) after receiving the newly updated information from the local server. Such performance evaluations, termed as rewards, could, for instance, be based on the ROUGE-2 score for automatic summarization tasks [41], and are represented by a random variable vector $\boldsymbol{X}_{t}=(X_{t,1},\ldots,X_{t,|S_{t}|})$. For simplification, we posit that $\boldsymbol{X}_{t}\in[0,1]^{|S_{t}|}$. Furthermore, for each LLM $k\in\mathcal{K}$ included in the action $S_{t}$, its associated cost $y_{t,k}$ is observable at round $t$. A detailed definition of this cost will be provided later.
Note that we demonstrate a fully online learning process here without any prior information for a clear description. In real-world scenarios, online learning can often be accelerated by incorporating offline pre-training. For example, for mathematical queries, narrowing the action space by offering fine-tuned LLMs for specific mathematical tasks, rather than selecting from a broad set, can significantly streamline the online selection process.
Versatile Reward Models. Let $\boldsymbol{\mu}=(\mu_{1},...,\mu_{K})$ represent the “initially unknown” mean vector of outcomes for each LLM with $\mu_{k}=\mathbb{E}[X_{t,k}]$. We consider versatile reward models for different combinations of LLMs in multi-LLM tasks below.
@@
  • All In Combination (AIC): $r(S;\boldsymbol{\mu})=\prod_{k\in S}\mu_{k}.$ This reward model is exemplified in development tasks, where each LLM is responsible for developing sub-modules simultaneously. The key aspect of this reward model is that the failure of any LLM leads to the failure of the entire development task, thus ensuring the success of the whole collaborative work.
For more discussions (e.g., specify LLMs under the SUC and AIC reward models), please refer to Section C.1. Compared to previous works [18, 13, 47, 73], which primarily focus on selecting a single LLM to satisfy the user, our versatile reward model accounts for the diversity of tasks, addressing scenarios where multi-LLMs collaborate to complete tasks, besides the distinction being our focus on an online learning setting.
Partial LLM Feedback. In the process of selecting and querying LLM in action $S_{t}$, not all LLMs may actually be queried in the muti-LLM tasks. Consequently, the local server often only observes outcomes from a subset of the selected LLMs. Formally, the local server observes feedback from the LLM subset $\mathcal{F}_{t}$ in action $S_{t}$, where the subset’s cardinality is denoted as $F_{t}$. If all LLMs in $S_{t}$ are queried, then $F_{t}=|S_{t}|$, which is worst-case scenario under the AWC reward model.
Statistically-Based Cost Model. Drawing inspiration from [1], which charges based on statistical computational resource utilization (with 1 CU supporting an average of 10 interactions), [3], which offers unlimited token pricing without the consideration of token limits, and [34], which bills based on exclusive resource groups, we propose a statistically-based cost model. Specifically, for any query $q$ from a set $\mathcal{Q}$, LLM $k\in\mathcal{K}$ processes the query using a deterministic number of input tokens, $l^{\text{in}}_{k}(q)$, and generates a random number of output tokens, denoted as $l^{\text{out}}_{k}(q)\sim D^{\text{out}}_{k}(q)$, representing the probabilistic nature of LLMs [70, 17]. Given a distribution $D_{q}$ over $\mathcal{Q}$, a user selects a query $q_{t}\sim D_{q}$ in each round, and the cost for LLM $k$ is $y_{t,k}=(l^{\text{in}}_{k}(q_{t})+l^{\text{out}}_{k}(q_{t}))C_{k}$, where $C_{k}$ is the cost per token. The goal is to estimate the unknown expected cost $c_{k}=\mathbb{E}[y_{t,k}]$.
Budget Violation Consideration. The total cost of executing action $S_{t}$ on the utilized LLM subset $\mathcal{F}_{t}$, selected by the scheduling cloud at round $t$, is given by $\sum_{k\in\mathcal{F}_{t}}y_{t,k}$. Furthermore, there is a predetermined budget guarantee threshold $\rho>0$ exists for the combinatorial set of LLM, requiring the cumulative cost of the selected action to remain below this threshold $\rho$ for the long term. This mechanism allows organizations, including enterprises and governmental bodies, to manage LLM usage efficiently and align with budgetary constraints. To assess compliance with this budgetary constraint, the concept of constraint violation [8] is introduced as follows, with $[x]^{+}:=\max(x,0)$:
$V(T)=\left[\frac{1}{T}\sum_{t=1}^{T}\sum_{k\in\mathcal{F}_{t}}y_{t,k}-\rho\right]^{+},$ (1)
Although it is conceivable to occasionally exceed this cost constraint during the online learning process, significant overruns are not permissible, thus necessitating the violation metric. A violation rate decreasing as $\tilde{O}\left(T^{-\gamma}\right)$ signifies effective constraint adherence, indicating that $V(T)$ diminishes at this rate, where $\gamma>0$ illustrates the rate of reduction in violations with $T$ [12].
@@ 4 Algorithm Design
Figure 3: Design of C2MAB-V workflow, with detailed process descriptions provided on the left main text.
We present the design of the online C2MAB-V framework, as depicted in Algorithm 1. Our method addresses the challenge of estimating unknown rewards and costs associated with using corresponding LLMs. As shown in Fig. 3, given the NP-hardness complexity of selecting a combination of LLMs under cost constraints, we transform the original integer problem into a continuous space. Utilizing the limited resources of a local server, we solve this relaxed optimization problem, which can also ease the computational load on the scheduling cloud when supporting multiple local-server query requests. A scheduling cloud with multiple LLMs coordinates and selects based on continuous data transmitted from the local server. Meanwhile, the local server collects user feedback to enhance the online evaluations of LLMs. Additionally, the cloud does not have access to the original sensitive user data, enhancing future privacy protection prospects. Next, we describe the processes undertaken by the local server and the scheduling cloud.
@@
1:Set of all LLMs $\mathcal{K}$, cost constraint $\rho$, probability parameter $\delta\in(0,1]$.
@@
4:     Predict reward of $k$-th LLM $\bar{\mu}_{t,k}=\min\{\hat{\mu}_{t,k}+\alpha_{\mu}\rho_{t,\mu_{k}},1\}$ based on “optimistic” strategy;
5:     Estimate cost of $k$-th LLM $\underaccent{\bar}{c}_{t,k}=\max\{\hat{c}_{t,k}-\alpha_{c}\rho_{t,c_{k}},0\}$ based on “pessimistic” strategy;
@@
8:     Observe the corresponding reward of LLMs and the cost of LLMs in for $k\in\mathcal{F}_{t}$ ;
@@ 4.1 Procedures by Local Server
Confidence Bound for Reward and Cost. To avoid the limitations of a greedy LLM selection strategy, i.e., $\operatorname*{arg\,max}_{k\in\mathcal{K}}$ $\hat{\mu}_{t,k}$ with $\hat{\mu}_{t,k}$ denoting the empirical estimate of the true mean of LLM $k$, which might result in the risk of overlooking superior LLM options, we implement the confidence bound (CB) method [39, 43]. This “optimistic” strategy promotes exploration among various LLM alternatives, thereby reducing the likelihood of consistently choosing sub-optimal options without considering potentially better LLMs. We define the confidence radius of CB as $\rho_{t,\mu_{k}}=\sqrt{\ln\left(\frac{2\pi^{2}Kt^{3}}{3\delta}\right)/2T_{t,\mu_{k}}}$, which quantifies the exploration potential for the reward of LLM $k$ at round $t$, with $\delta$ in the range $(0,1]$. The term $T_{t,\mu_{k}}$ represents the number of times LLM $k$ has been selected in action $S_{t}$. Accordingly, the adjusted reward prediction for LLM $k$ is defined as $\bar{\mu}_{t,k}=\min\{\hat{\mu}_{t,k}+\alpha_{\mu}\rho_{t,\mu_{k}},1\}$, with $\alpha_{\mu}$ denoted as a positive control parameter. Furthermore, we take into account the cost associated with each LLM. In light of the uncertainty introduced by complex queries, we adopt a cautious approach based on “pessimistic” strategy to ensure adherence to the cost budget. Specifically, the adjusted cost estimate for LLM $k$ at round $t$, $\underaccent{\bar}{c}_{t,k}$, is determined by $\underaccent{\bar}{c}_{t,k}=\max\{\hat{c}_{t,k}-\alpha_{c}\rho_{t,c_{k}},0\}$. Here, $\hat{c}_{t,k}$ is the empirical cost, $\alpha_{c}$ is a positive adjustment parameter, and confidence radiu $\rho_{t,c_{k}}=\sqrt{\ln\left(\frac{2\pi^{2}Kt^{3}}{3\delta}\right)/2T_{t,c_{k}}}$, with $T_{t,c_{k}}$ denoting the count of LLM $k$ selection up to round $t$. Using this CB approach, we have the following lemma.
@@
Please refer to Appendix D.1 for the proof. Lemma 1 underscores the high-probability events wherein the empirical estimates for both reward and cost closely align with their true means of LLMs.
@@
    The common greedy algorithms, apt for such constrained continuous problems, can efficiently select $\tilde{z}_{k}$ values that optimize $1-\bar{\mu}_{k}\tilde{z}_{k}$ within constraints [7]. Subsequently, this optimization problem can be easily and effectively solved.
@@
The commonalities in the above three types of reward forms are in Appendix C.2. We will further describe the selection of multiple LLMs based on the continuous variable $\tilde{\boldsymbol{Z}}_{t}=\{\tilde{z}_{t,1},\tilde{z}_{t,2},\cdots,\tilde{z}_{t,K}\}\in[0,1]^{K}$ at round $t$ in Section 4.2 on Procedures by Scheduling Cloud.
@@
$\hat{\mu}_{t+1,k}=\cfrac{T_{t,\mu_{k}}\hat{\mu}_{t,k}+X_{t,k}}{T_{t+1,\mu_{k}}},\quad\hat{c}_{t+1,k}=\cfrac{T_{t,c_{k}}\hat{c}_{t,k}+y_{t,k}}{T_{t+1,c_{k}}},\quad k\in\mathcal{F}_{t}.$ (6)
Note that for an action $S_{t}$, the local server only monitors the performance of the LLMs that are actually used in $S_{t}$, as not all LLMs are utilized for every type of task. For example, in the case of an AWC task type, if one LLM provides a satisfactory answer, the remaining LLMs are not utilized. In the AWC scenario, we use a conservative approach by ensuring that the budget threshold is met even if all selected LLMs $S_{t}$ are used, representing a “cautious” economic strategy.